Suppose the supply and demand curves for a particular product are given by: QS =-20+2P QD =100-2P. where QS and QD are quantities in units and P is the price per unit.

1
Expert's answer
2021-10-05T09:21:12-0400

"QS=QD\\-20+2P=100-2P\\4P=120\\P=30\\Q=40"